OpenStudy (heart_offire):

x = 3y - 6 2x - 4y = 8 Solve the system of equations using substitution.

rebeccaxhawaii (rebeccaxhawaii):

okay so x is 3y-6 so plug it into the second equation so 2(3y-6)-4y=8 then solve for y

OpenStudy (heart_offire):

@rebeccaxhawaii is y 2x-4y or..

rebeccaxhawaii (rebeccaxhawaii):

2(3y-6)-4y=8 solve for y

OpenStudy (heart_offire):

i distributed and got 6y - 16 - 4y =8

jimthompson5910 (jim_thompson5910):

how did you get 16?

OpenStudy (heart_offire):

6 x 2=12

OpenStudy (heart_offire):

i told you i distributed

jimthompson5910 (jim_thompson5910):

so it should be `6y - 12 - 4y =8 ` instead of `6y - 16 - 4y =8 ` (I changed the -16 to -12)

jimthompson5910 (jim_thompson5910):

what comes next after `6y - 12 - 4y =8` ?

OpenStudy (heart_offire):

umm do i solve it like aregular equation....subtract 12 from both sides

jimthompson5910 (jim_thompson5910):

what is 6y-4y equal to?

OpenStudy (heart_offire):

2y

jimthompson5910 (jim_thompson5910):

so we now have 2y-12 = 8 now add 12 to both sides to undo the -12

OpenStudy (heart_offire):

y=10

jimthompson5910 (jim_thompson5910):

y = 10 is correct

jimthompson5910 (jim_thompson5910):

then you'll go back to `x = 3y - 6` and plug in y = 10 to find x

OpenStudy (heart_offire):

x=24

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (heart_offire):

4x + 3y = 6 -4x + 2y = 14 Solve the system of equations. A) x = 1 2 , y = 3 B) x = 3, y = 1 2 C) x = 4, y = - 3 2 D) x = - 3 2 , y = 4

OpenStudy (heart_offire):

the numbers thats on top of each other are fractions